# Ergodic Risk Measures: Towards a Risk-Aware Foundation for Continual Reinforcement Learning

### v1

@@ 4 Continual Risk-Aware Reinforcement Learning
In this section, we present our primary contribution: the first formal theoretical treatment of risk-aware learning and decision-making in the continual setting. In particular, in Section 4.1, we show that both of the existing risk measure-based approaches that are used in non-continual risk-aware RL are incompatible with continual RL. Then, in Section 4.2, we build on this insight to propose a new class of ergodic risk measures, along with a corresponding RL objective, that are both compatible with continual RL. In Section 5, we leverage these results to showcase a case study in which we optimize an ergodic risk measure in a continual learning setting.
@@ 4.1 Existing Risk Measures and Continual Reinforcement Learning
One of the defining aspects of continual RL is the stability-plasticity dilemma, through which an agent must carefully balance the degree to which newly acquired information affects its behaviour relative to previously learned knowledge. Accordingly, in the risk-aware continual setting, we argue that this same dilemma should be reflected in how the agent assesses risk. In particular, we argue that any risk measure used in the continual setting should have some non-zero level of plasticity. That is, the risk evaluation at a given time step should depend only on the recent history leading up to that time step, rather than the entire history. However, in this section, we show that the static and nested risk measures used in non-continual risk-aware RL are not capable of such adaptability. To this end, we begin by formally defining two interpretations of plasticity as it relates to risk measures:
@@
(Fixed Plasticity) Let $(\Omega,\mathcal{F},\mathbb{P})$ denote a probability space with filtration $(\mathcal{F}_{t})_{t=0}^{N}$, where $N\in\mathbb{N}\cup\{\infty\}$, and let $\{\rho_{t}\}_{t=0}^{N}$ denote a sequence of conditional risk measures, such that $\rho_{t}:L^{\infty}(\mathcal{F}_{N})\to L^{\infty}(\mathcal{F}_{t})$, where $L^{\infty}(\mathcal{F}_{u})$ denotes the space of essentially bounded, $\mathcal{F}_{u}$-measurable random variables. The sequence $\{\rho_{t}\}$ is said to satisfy the fixed plasticity property if there exists a fixed, finite, non-zero horizon length, $m\in\mathbb{N},\;m<<N$, such that, for all $t\geq m$ and $X\in L^{\infty}(\mathcal{F}_{N})$, we have that: $\rho_{t}(X)\in L^{\infty}(\mathcal{G}_{t-m+1:t})$, where $\mathcal{G}_{t-m+1:t}$ denotes the $\sigma$-algebra generated by the information from the last $m$ time steps up to time $t$. That is, the risk evaluation at time $t$ depends only on the most recent $m$-step history.
Definition 4.2.
(Asymptotic Plasticity) Let $(\Omega,\mathcal{F},\mathbb{P})$ denote a probability space with filtration $(\mathcal{F}_{t})_{t=0}^{\infty}$, and let $\{\rho_{t}\}_{t=0}^{\infty}$ denote a sequence of conditional risk measures, such that $\rho_{t}:L^{\infty}(\mathcal{F}_{\infty})\to L^{\infty}(\mathcal{F}_{t})$. The sequence $\{\rho_{t}\}$ is said to satisfy the asymptotic plasticity property if there exists a finite time step, $n>>0$, such that, for all $X\in L^{\infty}(\mathcal{F}_{\infty})$, and some $\hat{\rho}_{t}(X)\in L^{\infty}(\mathcal{G}_{n+1:t})$, where $\mathcal{G}_{n+1:t}$ denotes the $\sigma$-algebra generated by the information from time step $n+1$ up to time step $t$, we have that: $\lim_{t\to\infty}\big\|\rho_{t}(X)-\hat{\rho}_{t}(X)\big\|_{\infty}=0$. That is, once sufficient time has passed, the risk evaluation effectively ceases to depend on any history that occurs prior to and including time step $n$.
In essence, the above definitions allow us to encode the notion of plasticity into risk measure theory. The first notion of plasticity (Definition 4.1) refers to fixed plasticity, which requires that risk evaluations depend only on the most recent $m$-sized window of history. Alternatively, the second notion of plasticity (Definition 4.2) refers to asymptotic plasticity, which instead requires that, over time, the influence of any past history vanishes, such that the risk evaluations effectively depend only on the more recent history. We note that (non-nested) Markov risk measures (see Definition A.7) can be viewed as satisfying the fixed plasticity definition with a window size of $m=1$.
As such, to be consistent with the stability-plasticity dilemma, we would expect that an appropriate risk measure in the continual setting is able to satisfy either the fixed plasticity property or the asymptotic plasticity property. However, we now show that the risk measures used in the static and nested approaches do not satisfy either of these properties:
@@
Let $\rho:L^{\infty}(\mathcal{F}_{j})\to\mathbb{R}$ denote a static risk measure (Definition A.2) defined for some time step, $j$. The static risk measure, $\rho$, does not satisfy the fixed plasticity property (Definition 4.1) or the asymptotic plasticity property (Definition 4.2).
@@
Definitions 4.1 and 4.2 require a sequence of conditional risk measures, $\{\rho_{t}\}_{t=0}^{N}$, such that $\rho_{t}:L^{\infty}(\mathcal{F}_{N})\to L^{\infty}(\mathcal{F}_{t})$, where $N\in\mathbb{N}\cup\{\infty\}$ for fixed plasticity (Definition 4.1) and $N=\infty$ for asymptotic plasticity (Definition 4.2). Conversely, a static risk measure is a single mapping corresponding to a single point in time. That is, it is not time-indexed, and therefore cannot adapt as new information arrives. Hence, it does not satisfy either definition of plasticity. ∎
Lemma 4.4.
Let $\rho_{t}:L^{\infty}(\mathcal{F}_{t+1})\to L^{\infty}(\sigma(S_{t},A_{t}))$ denote a (one-step) conditional Markov risk measure for some $S_{t}\in\mathcal{S},A_{t}\in\mathcal{A}$, and let $\rho_{0:N}$ denote a nested Markov risk measure over a time horizon, $N$, such that: $\rho_{0:N}(X)\doteq\rho_{0}\big(\rho_{1}(\cdots\rho_{N-1}(X)\cdots)\big)$ (see Definitions A.6 and A.7). The nested Markov risk measure, $\rho_{0:N}$, does not satisfy the fixed plasticity property (Definition 4.1) or the asymptotic plasticity property (Definition 4.2).
@@
Consider the nested risk measure structure: $\rho_{0:N}(X)=\rho_{0}\big(\rho_{1}(\rho_{2}(\cdots\rho_{N-1}(X)\cdots))\big)$. By the recursive structure, $\rho_{0}$ depends on $\rho_{1}(\cdot)$, which in turn depends on $\rho_{2}(\cdot)$, and so forth, until $\rho_{N-1}(X)$. This creates a dependency chain where: $\rho_{N-1}(X)$ depends on $(S_{N-1},A_{N-1})$; $\rho_{N-2}(\rho_{N-1}(X))$ depends on $(S_{N-2},A_{N-2})$ and, by the nested structure, $(S_{N-1},A_{N-1})$; and so forth until we have that $\rho_{0}(\rho_{1}(\cdots))$ depends on $(S_{0},A_{0},\ldots,S_{N-1},A_{N-1})$. As such, the risk evaluation depends on the entire history, thereby contradicting both definitions of plasticity. ∎
Altogether, Lemmas 4.3 and 4.4 show that the risk measure-based approaches used in non-continual risk-aware RL are incompatible with the continual setting.
4.2 Ergodic Risk Measures for Continual Reinforcement Learning
In the previous section, we showed that the risk measure-based approaches that are used for non-continual risk-aware RL (i.e., static and nested) are not compatible with the continual setting, and in particular, the stability-plasticity dilemma. Conversely, in this section, we propose a new class of ergodic risk measures, along with a corresponding RL objective, that are both compatible with the continual setting. To this end, having ruled out static and nested risk measures, let us begin by first considering what properties we would want a risk measure in the continual setting to satisfy:
  • (Non-Nested) Dynamic: We would indeed want a dynamic risk measure (i.e., a sequence of conditional risk measures) that can capture evolving risk preferences over time.
  • Coherent: The change from the non-continual to the continual setting does not affect this property. As such, while it is not strictly necessary to satisfy this property in the continual setting, it may still be beneficial to do so from a pure risk evaluation perspective.
  • Time-Consistent: The abstract notion of time consistency is, in itself, not necessarily incompatible with the continual setting; however, the way it is currently defined in the non-continual setting requires that a risk measure be time consistent for the entire history, which does go against the stability-plasticity dilemma (we will revisit this point below).
  • Plasticity: As discussed in Section 4.1, we would want the risk measure to satisfy either the fixed or asymptotic plasticity properties (see Definitions 4.1 and 4.2). We note that a (non-nested) Markov risk measure satisfies the fixed plasticity property.
Hence, in the continual setting, we would want a (non-nested) dynamic, possibly-coherent risk measure that satisfies one of the two plasticity properties. However, as mentioned above, there still remains the question of time consistency. In particular, the formal definition of time consistency in the non-continual setting (see Definition A.5) is clearly incompatible with the stability-plasticity dilemma (i.e., an agent should have the flexibility to change its risk preference over time). However, we can still define a weaker notion of time consistency that is compatible with the continual setting:
Definition 4.5.
(Local Time Consistency) Let $(\Omega,\mathcal{F},\mathbb{P})$ denote a probability space with filtration $(\mathcal{F}_{t})_{t=0}^{\infty}$, and let $\{\rho_{t}\}_{t=0}^{\infty}$ denote a sequence of conditional risk measures, such that $\rho_{t}:L^{\infty}(\mathcal{F}_{\infty})\to L^{\infty}(\mathcal{F}_{t})$. The sequence $\{\rho_{t}\}$ is said to satisfy the local time consistency property if there exist a time step, $n\geq 0$, and a possibly-infinite horizon length, $m\in\mathbb{N}\cup\{\infty\}$, such that, for all $n\leq t<n+m$ and all $X,X^{\prime}\in L^{\infty}(\mathcal{F}_{\infty})$, we have that: $\rho_{t+1}(X)\leq\rho_{t+1}(X^{\prime})\ \implies\ \rho_{t}(X)\leq\rho_{t}(X^{\prime})$. That is, time consistency holds within some subset of the time-horizon. Note that when $n=0$ and $m\to\infty$, this reduces to the standard definition of time consistency (Definition A.5).
In essence, the above definition for local time consistency only requires that time consistency holds for some subset of the history, rather than the entire history. We argue that such a notion of time consistency could be useful in a continual learning setting as it could provide some measure of stability. That is, while we want the agent to have the flexibility to change its risk preferences over time, it would likely be problematic if the agent changed its risk preferences at every time step.
As such, with the notion of time consistency now accounted for, we now have all the ingredients needed to define our proposed class of ergodic risk measures. In essence, an ergodic risk measure is a (non-nested) dynamic, possibly-coherent risk measure that satisfies the asymptotic plasticity and local time consistency properties. We provide a formal definition below:
Definition 4.6.
(Ergodic Risk Measure) Let $(\Omega,\mathcal{F},\mathbb{P})$ denote a probability space with filtration $(\mathcal{F}_{t})_{t=0}^{\infty}$, and let $\{\rho_{t}\}_{t=0}^{\infty}$ denote a sequence of conditional risk measures, such that $\rho_{t}:L^{\infty}(\mathcal{F}_{\infty})\to L^{\infty}(\mathcal{F}_{t})$. We call $\{\rho_{t}\}_{t=0}^{\infty}$ an ergodic risk measure if it satisfies the asymptotic plasticity property (Definition 4.2), and the local time consistency property (Definition 4.5).
Next, we can use our newly defined class of ergodic risk measures to motivate an appropriate RL objective for performing risk-aware learning and decision-making in the continual setting. To this end, let us consider the risk-aware analogue to the (risk-neutral) average-reward RL objective (1):
$\rho_{\pi_{t}}(s)\doteq\lim_{n\rightarrow{}\infty}\frac{1}{n}\sum_{t=t_{0}}^{n}\rho[R_{t}\mid S_{t_{0}}=s,A_{t_{0}:t-1}\sim\pi_{t}].$ (2)
That is, we want to optimize some risk measure, $\rho$, pertaining to the limiting per-step reward distribution induced when following a given (stationary) policy, $\pi_{t}$. However, the risk measure presented in Equation (2) is dependent on the initial conditions, which may not be ideal in the continual setting. As such, as with the average-reward objective (1), we can apply an appropriate ergodicity-like assumption that makes the risk-aware objective independent of the initial conditions. To this end, in this work, we utilize a unichain assumption for prediction (learning), and a communicating assumption for control (optimization):
Assumption 4.7 (Unichain Assumption for Prediction).
The Markov chain induced by the policy is unichain. That is, the induced Markov chain consists of a single recurrent class and a potentially-empty set of transient states.
Assumption 4.8 (Communicating Assumption for Control).
The MDP has a single communicating class. That is, each state in the MDP is accessible from every other state under some deterministic stationary policy.
Importantly, we can show that under the above ergodicity-like assumptions (or equivalent), the risk-aware objective (2) corresponds to an ergodic risk measure:
Theorem 4.9.
Given an appropriate ergodicity-like assumption, such as Assumption 4.7 or 4.8, and a stationary policy, $\pi_{t}$, the risk-aware objective (2) corresponds to an ergodic risk measure, as defined in Definition 4.6.
Proof.
Consider some arbitrary finite time step, $j>>t_{0}$. Under the ergodicity-like assumption, we can rewrite the risk-aware objective (2) as follows:
$\rho_{\pi_{t}}$ $\doteq\lim_{n\rightarrow{}\infty}\frac{1}{n}\sum_{t=t_{0}}^{n}\rho[R_{t}\mid S_{t}\sim\mu_{\pi_{t}},A_{t}\sim\pi_{t}]$ (3)
$=\lim_{n\rightarrow{}\infty}\frac{1}{n}\sum_{t=t_{0}}^{j}\rho[R_{t}\mid S_{t}\sim\mu_{\pi_{t}},A_{t}\sim\pi_{t}]+\lim_{n\rightarrow{}\infty}\frac{1}{n}\sum_{t=j+1}^{n}\rho[R_{t}\mid S_{t}\sim\mu_{\pi_{t}},A_{t}\sim\pi_{t}]$ (4)
$=0+\lim_{n\rightarrow{}\infty}\frac{1}{n}\sum_{t=j+1}^{n}\rho[R_{t}\mid S_{t}\sim\mu_{\pi_{t}},A_{t}\sim\pi_{t}]$ (5)
$=\rho_{\pi_{t}},$ (6)
where the final equality is due to Birkhoff’s Ergodic Theorem [Birkhoff, 1931], which ensures that $\rho_{\pi_{t}}$ converges to a stationary value as $t\to\infty$, regardless of the initial conditions. Hence, there exists some finite time step, $j$, such that the risk-aware objective effectively ceases to depend on any history that occurs prior to that time step, thereby satisfying the asymptotic plasticity requirement. Similarly, the local time consistency property follows directly from Birkhoff’s Ergodic Theorem, such that the risk-aware objective (2) is time-consistent as $t\to\infty$. Finally, the risk-aware objective (2) clearly evaluates risk over the entire time horizon, thereby satisfying the definition of a dynamic risk measure. Hence, all requirements of Definition 4.6 are satisfied. This completes the proof. ∎
As such, Theorem 4.9 establishes that, under ergodicity-like assumptions, the risk-aware objective (2) corresponds to an ergodic risk measure, thereby making it compatible with the continual setting.
@@ 4.2 Ergodic Risk Measures for Continual Reinforcement Learning
We note that ergodic risk measures are also compatible with the (generic) average-reward setting, given that they are capable of capturing distributional characteristics pertaining to the long-run per-step reward distribution induced when following a given stationary policy.

### v2

@@ 4 Continual Risk-Aware Reinforcement Learning
What does it mean for an RL agent to be risk-aware in a continual setting? In this section, we seek to answer this question by conducting the first formal theoretical treatment of risk-awareness in a continual setting. In particular, in Section 4.1, we formalize the notion of risk-awareness in a continual setting, and propose a set of axioms that establish the conditions needed to enable an agent to be risk-aware in a continual setting. Then in Section 4.2, we show that both of the existing classes of risk measures used in non-continual risk-aware RL (i.e., static and nested) are incompatible with these axioms. Finally, in Section 4.3, we propose a new class of ergodic risk measures, along with a corresponding RL objective, and show that both are compatible with continual learning.
4.1 Risk-Awareness and Continual Reinforcement Learning
In this section, we formalize the notion of risk-awareness in a continual setting. More specifically, we propose a risk-aware generalization of the continual RL framework proposed in Abel et al. [2025], such that we formalize risk-aware continual RL as an ongoing exchange of observations and actions between an agent and an environment, such that the agent receives observations from the environment, processes those observations into risk-aware observations based on its risk attitude (or risk tolerance), then emits actions back to the environment based on the risk-aware observations. We then leverage this risk-aware framework to propose a set of axioms that establish the conditions needed to enable an agent to be risk-aware in a continual setting.
To this end, let us begin by considering the aforementioned framework for continual RL proposed in Abel et al. [2025] and outlined in Section 3.1 of this work. Under this framework, one possible interpretation of the agent’s goal (or objective) is as follows:
Proposition 4.1.
Under the continual RL framework proposed in Abel et al. [2025], the agent’s goal (or objective) can be interpreted as wanting to leverage past observations from the environment, $O_{a:b}$, to find and output a sequence of actions, $A_{c:d}$, that will result in a future sequence of observations from the environment that optimizes some scalar objective, $J$.
Less formally, the above proposition states that the agent’s goal is to optimize some scalar objective, $J$. From a mathematical perspective, this $J$ is some functional, $J:\mathcal{O}\to\mathbb{R}$, that assigns to each observation, $O_{t}\in\mathcal{O}$, a scalar value, such that the agent’s goal is to induce a sequence of observations from the environment that optimizes this scalar value.
Now, let us consider this objective in the context of a risk-aware agent. In particular, when one examines the RL literature as it relates to risk-aware decision-making (e.g. see Section 2.2), there are two universal functions that a risk-aware agent needs to perform. First, the agent needs to be able to gauge or assess the risk of a given sequence of observations from the environment with respect to a given risk attitude or risk tolerance. Then, based on this risk assessment, the agent needs to find and output a sequence of actions that will result in a future sequence of observations whose risk assessment is optimal with respect to the agent’s risk attitude. Importantly, if we take the output of the agent’s risk assessment to be a scalar value, then the risk assessment process itself can be viewed as the mapping $\mathcal{O}\times\mathcal{T}\to\mathbb{R}$, where $\mathcal{T}$ denotes a set that captures the agent’s risk attitude. Accordingly, if we interpret $\mathcal{O}\times\mathcal{T}$ as a risk-aware observation, then the process of optimizing this observation (i.e., the assessed risk) can be viewed as a risk-aware generalization of the optimization process described in Proposition 4.1, where, by definition, $J$ corresponds to a risk measure, $\rho$ (see Section 2.2). This is formalized as Proposition 4.2 below:
Proposition 4.2.
If the output of the agent’s risk assessment is a scalar, then this scalar corresponds to a risk measure, and the act of optimizing this risk measure can be viewed as a risk-aware generalization of the optimization process described in Proposition 4.1, such that the agent receives observations from the environment, processes those observations into risk-aware observations based on its risk attitude, then emits actions back to the environment based on the risk-aware observations.
As such, we have essentially derived a risk-aware generalization of the continual RL framework proposed in Abel et al. [2025] (see Figure 1b)). Accordingly, we can also generalize the definition of plasticity (1) for a risk-aware agent as follows:
$\mathfrak{R}_{\begin{subarray}{c}a:b\\ c:d\end{subarray}}\doteq\mathbb{I}(O_{a:b}\times\tau_{a:b}\to A_{c:d}),$ (2)
where $O_{a:b}\times\tau_{a:b}$ denotes a sequence of risk-aware observations with respect to the agent’s risk attitude, $\tau_{{}_{t}}\in\mathcal{T}$, and $\mathfrak{R}$ denotes the risk-aware plasticity.
Before proceeding, we want to emphasize the point made about $\rho$ being a generalization of $J$. In particular, if we take $\rho(X,\tau)=\mathbb{E}[X]$ (either by explicitly setting $\rho=\mathbb{E}$, or based on a specific choice of $\tau\in\mathcal{T}$ that results in $\rho(X,\tau)=\mathbb{E}[X]$), then we recover the risk-neutral objective, $J$. Accordingly, we note that any result or discussion point that is made with reference to a risk-aware objective, $\rho$, in this work is also applicable to any risk-neutral objective, $J$.
4.1.1 Axioms for Risk-Awareness in the Continual Setting
Having formalized the notion of risk-awareness in the context of continual RL, we now seek to establish the minimum requirements needed for the agent to be able to perform the aforementioned risk assessment and optimization in a continual learning setting. To this end, we argue that any limitations related to the risk assessment and subsequent optimization should be direct consequences of the limitations of the agent itself. In particular, we focus on two well-argued limitations (or requirements) of continual RL agents (e.g. see Abel et al. [2023], Kumar et al. [2025]), which are formalized as Assumption 4.3 below:
Assumption 4.3.
Any continual RL agent has a finite memory, and has a non-zero level of plasticity.
As such, based on the above limitations of the agent, we conclude this section by proposing the following axioms as the minimum requirements needed for risk-awareness in the continual setting:
Proposition 4.4.
Under Assumption 4.3, the following two axioms are necessary and sufficient conditions for risk-awareness in the continual setting: 1. Feasibility Axiom: The agent’s risk assessment must be computable based on the agent’s finite memory. That is, the risk assessment must be computable based on observations collected over a finite time interval $[a:b]$ with respect to the agent’s risk attitude during that same time interval. 2. Plasticity Axiom: $\mathbb{P}(\mathfrak{R}>0\text{ infinitely often})=1$. In words, this axiom states that the agent’s risk assessment and subsequent behaviour should be influenced by changes in its risk attitude as well as new observations from the environment infinitely often.
@@ 4.2 Existing Risk Measures and Continual Reinforcement Learning
In the previous section, we formalized the notion of risk-awareness in the continual setting, and showed that risk measures are a natural fit as an objective in this setting. But how exactly should we go about designing a risk measure? That is, how can we design a risk measure that satisfies the axioms in Proposition 4.4, and is also a sensible and ‘good’ measure of risk?
A natural first step towards answering this question could be to examine the existing classes of risk measures used in non-continual risk-aware RL (see Section 3.2.1) and see if we can leverage them in the continual setting. To this end, upon examining the existing classes of risk measures used in non-continual risk-aware RL (i.e., static and nested), we find that they are incompatible with the axioms established in Proposition 4.4. This is formally shown below:
Lemma 4.5.
Let $\rho_{N}$ denote a static risk measure (as per Definition A.2) defined for some time step, NN. The static risk measure, $\rho_{N}$, does not satisfy both axioms of Proposition 4.4.
Proof.
As $N\to\infty$, this violates the Feasibility axiom as the computation of the risk would require the agent to have infinite memory. If $N<\infty$, then any observation that occurs after time step $N$ would not influence the risk assessment, thereby violating the Plasticity axiom. ∎
Lemma 4.6.
Let $\{\rho_{t}\}_{t=0}^{\infty}$ denote a nested risk measure and let $\rho_{n}$ denote the assessed risk at time step $n$, such that its value can be computed as: $\rho_{n}(X)\doteq\rho_{n}\big(\rho_{n+1}(\cdots)\big)$ (see Definition A.6). The nested risk measure, $\{\rho_{t}\}_{t=0}^{\infty}$, does not satisfy both axioms of Proposition 4.4.
Proof.
Consider the risk assessment at time $n$, $\rho_{n}(X)\doteq\rho_{n}\big(\rho_{n+1}(\cdots)\big)$. By definition, the validity of this risk calculation is contingent on time consistency holding for all $t\geq n$. However, this implies that future observations cannot influence how risky $X$ is relative to other outcomes. Hence, either time consistency is satisfied, which violates the Plasticity axiom, or time consistency is violated, which invalidates the risk calculation at time $n$, thereby violating the Feasibility axiom. ∎
4.3 Ergodic Risk Measures for Continual Reinforcement Learning
In the previous section, we showed that static and nested risk measures are not compatible with continual learning. Based on this insight, we now propose a new class of ergodic risk measures, along with a corresponding RL objective, and show that they are both compatible with continual learning. To this end, we begin by formally defining our proposed class of ergodic risk measures:
@@ 4.3 Ergodic Risk Measures for Continual Reinforcement Learning
(Ergodic Risk Measure) Let $\{\rho_{t}\}_{t=0}^{\infty}$ denote a sequence of conditional risk measures (i.e., a dynamic risk measure) defined on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ with filtration $(\mathcal{F}_{t})_{t=0}^{\infty}$, where $\mathcal{F}_{t}\subseteq\mathcal{F}$ denotes the $\sigma$-algebra representing the information collected up until time step $t\geq 0$, such that $\rho_{t}:L^{\infty}(\mathcal{F}_{\infty})\to L^{\infty}(\mathcal{F}_{t})$, where $L^{\infty}(\mathcal{F}_{u})$ denotes the space of essentially bounded, $\mathcal{F}_{u}$-measurable random variables.
Suppose that for any time step, $t>0$, there exists a finite time interval, $[n:t]$, such that for all $X\in L^{\infty}(\mathcal{F}_{\infty})$, and some $\hat{\rho}_{t}(X)\in L^{\infty}(\mathcal{G}_{n:t})$, where $\mathcal{G}_{n:t}$ denotes the $\sigma$-algebra generated by the information from time step $n$ up to time step $t$, we have that: $\rho_{t}(X)-\hat{\rho}_{t}(X)\approx 0.$ (3) That is, the risk assessment at any given time step, $t$, can be accurately computed based on a preceding, finite subset of the history, such that the risk assessment effectively ceases to depend on the history that occurs prior to the interval $[n:t]$. Furthermore, suppose that $\{\rho_{t}\}_{t=0}^{\infty}$ is designed in such a way that there exists some time step, $t$, such that time consistency (Definition A.5) does not hold. Then, we call $\{\rho_{t}\}_{t=0}^{\infty}$ an ergodic risk measure.
In essence, an ergodic risk measure is a (non-nested) dynamic risk measure, such that its one-step risk assessments, $\rho_{t}$, can be accurately computed using only observations collected over a rolling finite time interval, $[n:t]$. We note that the size of this interval can vary between different time steps. Figure 2 depicts a comparison between static, nested, and ergodic risk measures. In particular, the figure visualizes which observations are needed to accurately compute the risk at a given time step.
Figure 2: A comparison between: a) static, b) nested, and c) ergodic risk measures in terms of which observations are needed to accurately compute the risk at a given time step. A shaded box indicates that the observation (as per the y-axis value) is needed to accurately compute the risk at that time step (x-axis value). Note that the nested risk measure depicted in this figure is assumed to be a Markov risk measure (see Definition A.7).
We now show that ergodic risk measures satisfy both of the axioms from Proposition 4.4:
@@
Let $\{\rho_{t}\}_{t=0}^{\infty}$ denote an ergodic risk measure as defined in Definition 4.7. The ergodic risk measure, $\{\rho_{t}\}_{t=0}^{\infty}$, satisfies both axioms of Proposition 4.4.
@@
Since an ergodic risk measure is a dynamic risk measure such that time consistency (as defined in Definition A.5) does not hold, it satisfies the Plasticity axiom since the risk assessments and subsequent behaviour will be influenced by new observations and any changes in the agent’s risk attitude infinitely often. Similarly, by definition, the risk at any time $t$ can be computed using observations collected over a finite time interval with respect to the agent’s risk attitude during that same time interval, thereby satisfying the Feasibility axiom. This completes the proof. ∎
4.3.1 An Ergodic Reinforcement Learning Objective
Having established that ergodic risk measures are compatible with continual learning, we now motivate a corresponding RL objective that can be used to perform risk-aware learning and decision-making in the continual setting. To this end, we will utilize the average-reward MDP framework [Puterman, 1994] to realize the agent-environment interaction described in Section 3.1. Our choice of the average-reward MDP formulation in this work follows prior work, such as Sharma et al. [2022] and Kumar et al. [2025], which argue that the average-reward formulation’s emphasis on long-term performance is a natural fit for continual learning settings.
More formally, consider a finite MDP, $\mathcal{M}\doteq\langle\mathcal{S},\mathcal{A},\mathcal{R},p\rangle$, where $\mathcal{S}$ is a finite set of states, $\mathcal{A}$ is a finite set of actions, $\mathcal{R}\subset\mathbb{R}$ is a bounded set of rewards, and $p:\mathcal{S}\,\times\,\mathcal{A}\,\times\,\mathcal{R}\,\times\,\mathcal{S}\rightarrow{}[0,1]$ is a probabilistic transition function that describes the dynamics of the environment, such that at each discrete time step, $t=0,1,2,\ldots$, an agent chooses an action, $A_{t}\in\mathcal{A}$, based on its current state, $S_{t}\in\mathcal{S}$, and receives a reward, $R_{t+1}\in\mathcal{R}$, while transitioning to a (potentially) new state, $S_{t+1}$, such that $p(s^{\prime},r\mid s,a)=\mathbb{P}(S_{t+1}=s^{\prime},R_{t+1}=r\mid S_{t}=s,A_{t}=a)$.
Under this framework, the continual RL problem can be viewed as an infinite sequence of MDPs, $\{\mathcal{M}^{k}\}_{k=1}^{\infty}$, such that $\mathcal{M}^{k}\doteq\langle\mathcal{S}_{k},\mathcal{A}_{k},\mathcal{R}_{k},p_{k}\rangle$, where each $\mathcal{M}^{k}$ may differ in its state-space, action-space, reward function, and/or transition dynamics based on some indexing function, $\omega:\mathbb{N}\to\mathbb{N}$, such that $\omega(t)=i$ indicates that at time step $t$, the agent is in MDP $\mathcal{M}^{i}$.
The agent’s goal in this framework is to construct a sequence of stationary policies, $\{\pi^{k}\}_{k=1}^{\infty}$, that optimizes the long-run (or limiting) average-reward, $\bar{r}$, which is defined as follows for a given stationary policy being followed at time $t$, $\pi_{t}\doteq\pi^{\omega(t)}$:
$\bar{r}_{\pi_{t}}(s)\doteq\lim_{h\rightarrow{}\infty}\frac{1}{h}\sum_{n=n_{0}}^{h}\mathbb{E}[R_{n}\mid S_{n_{0}}=s,A_{n_{0}:n-1}\sim\pi^{\omega(n_{0})}],$ (4)
where $n_{0}=\min\{y:\pi^{\omega(y)}=\pi_{t}\}$. When working with average-reward MDPs, it is common to simplify the expression for the average-reward objective (4) into a more workable form by making certain ergodicity-like assumptions about the Markov chain induced when following policy $\pi_{t}$. To this end, a unichain assumption is typically used when doing prediction (learning) because it ensures the existence of a unique limiting distribution of states, $\mu_{\pi_{t}}(s)\doteq\lim_{n\rightarrow{}\infty}\mathbb{P}(S_{n}=s\mid A_{n_{0}:n-1}\sim\pi^{\omega(n_{0})})$, that is independent of initial conditions. Similarly, a communicating assumption is typically used for control (optimization) because it ensures the existence of a unique optimal average-reward, $\bar{r}*$, that is independent of initial conditions. Importantly, these ergodicity-like assumptions enable the agent’s objective to be expressed as a stable measure of long-term performance that eventually becomes independent of prior conditions (i.e., $\bar{r}_{\pi_{t}}(s)=\bar{r}_{\pi_{t}}$).
Hence, to perform risk-aware learning and decision-making in a continual setting, we can consider the risk-aware analogue to the (risk-neutral) average-reward RL objective (4):
$\rho_{\pi_{t}}(s)\doteq\lim_{h\rightarrow{}\infty}\frac{1}{h}\sum_{n=n_{0}}^{h}\rho_{\tau}[R_{n}\mid S_{n_{0}}=s,A_{n_{0}:n-1}\sim\pi^{\omega(n_{0})}].$ (5)
That is, we want to optimize some risk measure, $\rho$, pertaining to the limiting per-step reward distribution induced when following a given (stationary) policy, $\pi_{t}$. However, the risk measure presented in Equation (5) is dependent on the initial conditions, which is not tractable in the continual setting. As such, as with the average-reward objective (4), we can apply an appropriate ergodicity-like assumption that makes the risk-aware objective independent of prior conditions (i.e., $\rho_{\pi_{t}}(s)=\rho_{\pi_{t}}$):
Assumption 4.9 (Unichain Assumption for Prediction).
For a given MDP, $\mathcal{M}^{k}$, the Markov chain induced by the policy, $\pi^{k}$, is unichain. That is, the induced Markov chain consists of a single recurrent class and a potentially-empty set of transient states.
Assumption 4.10 (Communicating Assumption for Control).
For a given MDP, $\mathcal{M}^{k}$, the MDP has a single communicating class. That is, each state in the MDP is accessible from every other state in the MDP under some deterministic stationary policy.
For clarity, we note that the above assumptions need only apply independently for each MDP, $\mathcal{M}^{k}$, rather than holistically for the entire sequence of MDPs, $\{\mathcal{M}^{k}\}_{k=1}^{\infty}$. For example, Assumption 4.10 only requires that each state in a given MDP, $\mathcal{M}^{k}$, is accessible from every other state in $\mathcal{M}^{k}$, but not the states of other MDPs, $\mathcal{M}^{i\neq k}$.
We now show that, under the above ergodicity-like assumptions, the risk-aware objective (5) corresponds to an ergodic risk measure, thereby making it compatible with continual learning:
Theorem 4.11.
Given an appropriate ergodicity-like assumption, such as Assumption 4.9 or 4.10, and a stationary policy, $\pi_{t}$, the risk-aware objective (5) corresponds to an ergodic risk measure, as defined in Definition 4.7.
@@ 4.3.1 An Ergodic Reinforcement Learning Objective
We begin by noting that the risk-aware objective (5) clearly assesses risk over the entire time horizon and does not enforce time consistency (as per Definition A.5), thereby satisfying the definition of a dynamic, non time-consistent risk measure.
Next, we show that for any time step, $t$, we can accurately compute the risk based on observations collected over a finite time interval. To this end, consider an MDP, $\mathcal{M}^{k}$, through which an agent interacts with the environment for the possibly-infinite time interval $[a:b]$ (note that $a$ here is equivalent to $n_{0}$ in Equation (5)). If $[a:b]$ is finite, then the desired condition is automatically satisfied. If $[a:b]$ is infinite, then, under the ergodicity-like assumption, we have, by Birkhoff’s Ergodic Theorem [Birkhoff, 1931], that $\rho_{\pi_{t}}$ converges to a stationary value as $t\to\infty$. This implies that there exists a finite time step, $j$, such that for all $t\geq j$, $\rho_{\pi_{t}}$ converges to a stationary value. We therefore have two cases to check: 1) If $t\leq j$, then, since $j$ is finite, the risk can be computed based on observations collected over the finite time interval $[a:j]$, thereby satisfying the desired condition. 2) If $t>j$, then the convergence of $\rho_{\pi_{t}}$ to a stationary value as $t\to\infty$ implies that for any $t>j$, we can find a finite time interval, $[n:t]$ (where $n\geq j$), that can be used to accurately compute the risk assessment, thereby satisfying the desired condition.
As such, the definition of an ergodic risk measure (4.7) is satisfied. This completes the proof. ∎
We conclude this section by providing some remarks related to ergodic risk measures:
@@
Ergodic risk measures are also compatible with the generic, possibly non-continual average-reward setting. That is, ergodic risk measures provide a formalism for risk-aware objectives of the form (5), which had been studied previously in non-continual average-reward settings (e.g. Xia et al. [2023], Rojas and Lee [2025]), but never formalized in the context of risk measure theory.